Aspen Company estimates its manufacturing overhead to be $848,000 and its direct labor costs to be $530,000 for year 2. Aspen worked on three jobs for the year. Job 2-1, which was sold during year 2, had actual direct labor costs of $166,000. Job 2-2, which was completed, but not sold at the end of the year, had actual direct labor costs of $291,000. Job 2-3, which is still in work-in-process inventory, had actual direct labor costs of $116,000. Actual manufacturing overhead for year 2 was $915,000. Manufacturing overhead is applied on the basis of direct labor costs.

Required:

a. How much overhead was applied to each job in year 2?

b. What was the over- or underapplied manufacturing overhead for year 2?
Answer #1
Estimated manufacturing overhead 848000
Divide by Estimated Direct labor costs 530000
Predetermined overhead rate 160%
a
Job 2-1 265600 =166000*160%
Job 2-2 465600 =291000*160%
Job 2-3 185600 =116000*160%
b
Actual direct labor costs 915000
Less: Overhead applied 916800 =265600+465600+185600
Overapplied manufacturing overhead 1800